chuckthedonut:

The circumference of a circle is 26 5/7 ft. What is the approximate diameter of the circle? Use 22/7 for π. Enter your answer as a mixed number in simplest form in the box.

chuckthedonut:

@vocaloid can you help?

chuckthedonut:

@oliver69, @jhonyy9, @shadow can you help?

chuckthedonut:

Can someone help me?

jhonyy9:

formula for circumference of a circle is C = 2*pi*r C = d*pi where d is diameter what is 2r